b) Identify and trace the conic x2 −2xy+ y2 −3x+ 2y+ 3 = 0.

1
Expert's answer
2021-05-31T16:06:06-0400

General Cartesian form of the equation is:


"Ax^2+Bxy+Cy^2+Dx+Ey+F=0"

If discriminant is equal to zero, the equation represents a parabola.

"B^2-4AC=2^2-4=0"

Therefore it is a parabola.